\begin{thebibliography}{11}
\bibitem{HOOFT} G. 't Hooft, Nucl. Phys. {\bf 72} (1974) 461;
%\bibitem{WITTEN}
E. Witten, Nucl. Phys. {\bf B160} (1979) 57.

%\bibitem{TRENTO} Proceedings of the Workshop {\it Large $N_c$ QCD 2004}, 
%Trento, Italy, July 5-11 2004, eds. J. L. Goity, R. F. Lebed, A. Pich, 
%C. L. Schat and N. N. Scoccola (World Scientific, Singapore, 2005).

%\cite{Gervais:1983wq}
\bibitem{Gervais:1983wq}
  J.~L.~Gervais and B.~Sakita,
%%%   ``Large N QCD Baryon Dynamics: 
%%%     Exact Results From Its Relation To The Static
  %Strong Coupling Theory,''
  Phys.\ Rev.\ Lett.\  {\bf 52} (1984) 87;
  %%CITATION = PRLTA,52,87;%%
%\cite{Gervais:1984rc}
%%%\bibitem{Gervais:1984rc}
%%%  J.~L.~Gervais and B.~Sakita,
  %``Large-N baryonic soliton and quarks,''
  Phys.\ Rev.\ D {\bf 30} (1984) 1795.
  %%CITATION = PHRVA,D30,1795;%%



\bibitem{DM} R. Dashen and A. V. Manohar, Phys. Lett. {\bf B315} (1993) 425;
ibid {\bf B315} (1993) 438.

%\bibitem{DJM94}
%R.~Dashen, E.~Jenkins, and A.~V.~Manohar,
%Phys.~Rev.  {\bf D49} (1994) 4713.

\bibitem{DJM95}
R. Dashen, E. Jenkins, and A. V. Manohar,
Phys. Rev. {\bf D51} (1995) 3697.

%\cite{Jenkins:1998wy}
\bibitem{Jenkins:1998wy}
  E.~Jenkins,
  %``Large-N(c) baryons,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 48} (1998) 81
%  [arXiv:hep-ph/9803349].
  %%CITATION = HEP-PH 9803349;%%


%\bibitem{CGO94}
%C.~D.~Carone, H.~Georgi and S.~Osofsky,
%Phys.~Lett. {\bf B322} (1994) 227.\\
%M.~A.~Luty and J.~March-Russell,
%Nucl.~Phys. {\bf B426} (1994) 71.\\
%M.~A.~Luty, J.~March-Russell and  M.~White,
%Phys.~Rev.  {\bf D51} (1995)  2332.

%\bibitem{Jenk1}
%E.~Jenkins, Phys.~Lett.  {\bf B315} (1993) 441.

%\bibitem{JL95}
%E.~Jenkins and R.~F.~Lebed,
%Phys.~Rev.  {\bf D52} (1995) 282.

%\bibitem{DDJM96}
%J.~Dai, R.~Dashen, E.~Jenkins, and A.~V.~Manohar,
%Phys.~Rev.   {\bf D53} (1996) 273.

%\bibitem{CGKM} C. D. Carone, H. Georgi, L. Kaplan and D. Morin,
% Phys. Rev. {\bf D50} (1994) 5793.

\bibitem{Goi97}
J.~L. Goity, Phys. Lett.  {\bf B414} (1997) 140.

%\bibitem{PY1} D. Pirjol and T.~M. Yan,
% Phys. Rev.  {\bf D57}  (1998) 1449.

%\bibitem{PY2}  D. Pirjol and T.~M. Yan,  Phys. Rev. {\bf D57} (1998) 5434.

\bibitem{CCGL}
C.~E. Carlson, C.~D. Carone, J.~L. Goity and R.~F. Lebed,
% Phys. Lett. {\bf B438}  (1998) 327;
 Phys. Rev. {\bf D59}  (1999) 114008.


\bibitem{CaCa98}
C.~E.~Carlson and C.~D.~Carone,
% Phys.\ Lett.\  {\bf  B441} (1998) 363;
 Phys.\ Rev.\  {\bf  D58} (1998) 053005.

%\bibitem{BCCG} Z. A. Baccouche, C. K. Chow, T. D. Cohen and B. A. Gelman,
% Nucl.\ Phys.\  {\bf  A696}  (2001) 638.
 
%\bibitem{SCHAT} C. L. Schat, 
%\emph{Tempe 2002, Phenomenology of large N(c) QCD}, (2002) 189-198 [arXiv:hep-ph/0204044].

%\bibitem{Pirjol:2003ye}
%D.~Pirjol and C.~Schat, Phys. Rev. {\bf  D67} (2003) 096009.

%\bibitem{cohen1}
%T. D. Cohen, D. C. Dakin, A. Nellore and R. F. Lebed,
%Phys. Rev. {\bf D69} (2004) 056001.

\bibitem{SGS}
%C.~L. Schat, J.~L. Goity and N.~N. Scoccola,  Phys. Rev. Lett.
%{\bf 88} (2002) 102002;
J.~L.~Goity, C.~L.~Schat and N.~N.~Scoccola,
 Phys.\ Rev.\  {\bf  D66} (2002) 114014.
 

\bibitem{CC00}
C. E. Carlson and C. D. Carone,
Phys. Lett. {\bf B484} (2000) 260.

\bibitem{GSS}
J.~L. Goity,   C.~L.~Schat and N.~N.~Scoccola,
 Phys. Lett. {\bf B564} (2003) 83.

\bibitem{MS2} N. Matagne and Fl. Stancu, Phys. Lett {\bf B631} (2005) 7.




%\cite{Matagne:2006zf}
\bibitem{Matagne:2006zf}
  N.~Matagne and Fl.~Stancu,
  %``Masses of (70,l+) baryons in the 1/N(c) expansion,''
  Phys. Rev. {\bf D74} (2006) 034014
%  [arXiv:hep-ph/0604122].
  %%CITATION = HEP-PH 0604122;%%

%\bibitem{GOITY05} J. L. Goity, Proceedings of the Workshop {\it Large $N_c$ QCD 2004}, 
%(World Scientific, Singapore, 2005) 211 or [arXiv:hep-ph/0504121].
   
\bibitem{MS1} N. Matagne and Fl. Stancu,  Phys. Rev. {\bf D71} (2005) 014010.

%\bibitem{SS94} P. Stassart and Fl. Stancu, Z. Phys. {\bf A359} (1997) 321.


%\emph{Trento 2004, Large N(c) QCD}, (2004) 211-222 [arXiv:hep-ph/0504121].

\bibitem{HP} K. T. Hecht and S. C. Pang, J. Math. Phys. {\bf 10} (1969) 1571.

%\cite{Matagne:2006xx}
\bibitem{Matagne:2006xx}
  N.~Matagne and Fl.~Stancu,
  %``Matrix elements of SU(6) generators for baryons at arbitrary N(c),''
  Phys. Rev.  {\bf D73} (2006) 114025
%  [arXiv:hep-ph/0603032].
  %%CITATION = HEP-PH 0603032;%%



%\bibitem{MS3} N. Matagne and Fl. Stancu, 
%Phys.\ Rev.\ D {\bf 74}, 034014 (2006) [arXiv:hep-ph/0603032].

%\bibitem{PS} D. Pirjol and C. Schat,  [arXiv:hep-ph/0603127].

%\bibitem{CL} T. D. Cohen and R. F. Lebed, Phys. Rev. {\bf D70} (2004) 096015.


 
%\bibitem{cohen2} T. D. Cohen and  R. F. Lebed,
% Phys. Rev. {\bf D67}, 096008 (2003); Phys. Rev. Lett. {\bf 91}, 012001 (2003);
% Phys. Rev. {\bf D68}, 056003 (2003).


%\bibitem{book} Fl. Stancu, {\it Group Theory in Subnuclear Physics},
%Clarendon Press, Oxford, (1996) Ch. 4.

%\bibitem{ISOSC} Fl. Stancu and S. Pepin, Few-Body Systems {\bf 26} (2004) 113.

%\bibitem{GR} L. Ya. Glozman and D. O. Riska, Phys. Rept. {\bf 268} (1996) 263.

%\bibitem{BRINK} D. M. Brink and G. R. Satchler {\it Angular Momentum}
%Oxford University Press, 1968, Ch. IV.

%\bibitem{HECHT} K. T. Hecht, Nucl. Phys. {\bf 62} (1965) 1.


%\bibitem{DESWART} J. J. De Swart, Rev. Mod. Phys. {\bf 35} (1963) 916.

%\bibitem{ELLIOTT} J. P. Elliott, 
%Proc. Roy. Soc. (London) {\bf A245} (1958) 128; ibid. {\bf A245} (1958) 562.

%\cite{Matagne:2006dj}
\bibitem{Matagne:2006dj}
  N.~Matagne and Fl.~Stancu,
  %``New look at the (70,1-) baryon multiplet in the 1/N(c) expansion,''
  arXiv:hep-ph/0610099.
  %%CITATION = HEP-PH 0610099;%%


\bibitem{Yao}
  W.-M. Yao {\it et al.} [Particle Data Group], J. Phys. {\bf G33} (2006) 1.





\end{thebibliography}

\documentclass[final,numberedheadings,cmfonts]{aipproc}
\layoutstyle{6x9}
\usepackage{multirow}

\def\1{\mbox{l\hspace{-0.53em}1}}

\begin{document}

\title{The $[{\bf 70},1^-]$ baryon multiplet in the $1/N_c$ expansion revisited}

\classification{12.39.-x,11.15.Pg,11.30.Hv}
\keywords      {baryon spectroscopy, large $N_c$ QCD}

\author{N. Matagne and Fl. Stancu}{address={University of Li\`ege, Physics Department,\\
Institute of Physics, B5, \\ 
Sart Tilman, B-4000 Li\`ege 1, Belgium\\ 
E-mail: nmatagne@ulg.ac.be,
fstancu@ulg.ac.be}}

\begin{abstract}

The mass splittings of the baryons belonging to the $[{\bf 70},1^-]$-plet
are derived by using a simple group theoretical approach to the matrix elements 
of the mass formula. 
The basic conclusion is that
the first order correction to the baryon masses is of order 
$1/N_c$ instead of order $N^0_c$, as previously found. 
The conceptual difference between the ground state and the excited states is therefore removed.
\end{abstract}

\maketitle

{\it Introduction.}
The $1/N_c$ expansion of QCD  \cite{HOOFT}
is a powerful theoretical tool which allows to systematically analyze
baryon properties.
The success of the method stems from the discovery that the ground state baryons have 
an exact contracted SU(2$N_f$) symmetry 
when $N_c \rightarrow \infty $  \cite{Gervais:1983wq,DM},  $N_f$
being the number of flavors.
A considerable amount of work has been devoted to the ground state
baryons,
summarized in several review papers as, {\it e. g.}, 
\cite{DJM95,Jenkins:1998wy}.
For $N_c \rightarrow \infty $ the ground state baryons are degenerate. 
For large $N_c$ the mass splitting starts at order $1/N_c$.
The applicability of the approach to excited states is a subject of
current investigation. The  experimental facts indicate a small breaking
of SU($2N_f$) which make the $1/N_c$ studies of excited states plausible.
When the SU($N_f$) symmetry is exact, the baryon mass operator is 
a linear combination of terms 
\begin{equation}
\label{massoperator}
M  = \sum_{i} c_i O_i,
\end{equation} 
with the operators $O_i$  having the general form
\begin{equation}\label{OLFS}
O_i = \frac{1}{N^{n-1}_c} O^{(k)}_{\ell} \cdot O^{(k)}_{SF},
\end{equation}
where  $O^{(k)}_{\ell}$ is a $k$-rank tensor in SO(3) and  $O^{(k)}_{SF}$
a $k$-rank tensor in SU(2), but invariant in SU($N_f$).
The latter is expressed in terms of  SU($N_f$) generators. 
For the ground state one has $k$ = 0. The first factor gives the 
order $\mathcal{O}(1/N_c)$ of the operator in the series expansion.
The lower index $i$ represents a specific combination of generators, see
Table \ref{operators}. 
In  Eq. (\ref{massoperator}), each $O_i$ is multiplied 
by an unknown coefficient $c_i$ which is a reduced matrix element.
All these coefficients encode the QCD dynamics and are obtained from a fit to
the existing data. 
Additional terms are needed if SU($N_f$) is broken.

\vspace{0.3cm}


{\it Excited states.}
The excited states can be grouped into excitation bands 
with N = 1, 2, 3, etc. units of excitation energy.
Among these, the states belonging to the N = 1 band,
described by the $[{\bf 70},1^-]$-plet, 
have been most extensively studied,
either for $N_f$ = 2, see {\it e.g.} 
\cite{Goi97,CCGL,CaCa98}
or for  $N_f$ = 3 \cite{SGS}. 
The conclusion was that the splitting starts 
at order $N^0_c$.

The method has  been applied to the N = 2 band multiplets 
$[{\bf 56'},0^+]$ for $N_f$ = 2  \cite{CC00},  $[{\bf 56},2^+]$ 
for $N_f = 3$ \cite{GSS}  and  $[{\bf 70},\ell^+]$   for 
$N_f = 2$ \cite{MS2} and $N_f = 3$ \cite{Matagne:2006zf}
and  also to   $N_f = 3$
baryons \cite{MS1} of the  $[{\bf 56},4^+]$-plet
(N = 4 band).

The excited states belonging to  $[{\bf 56},\ell]$ multiplets are rather
simple and can be studied by analogy to the ground state. 
Naturally the splitting starts at order $1/N_c$ \ \cite{GSS,MS1}.

\vspace{0.2cm}
\begin{table}[ph]
\caption{List of operators and  coefficients of the mass operator
(\ref{massoperator}) for $[{\bf 70},1^-]$.}
\label{operators}
{\tiny
\renewcommand{\arraystretch}{2} 
\begin{tabular}{llrrrrr}
\hline
Operator &  & Fit 1 (MeV) &  Fit 2 (MeV) & Fit 3 (Mev) & Fit 4 (MeV) & Fit 5 (MeV) \\
\hline
$O_1 = N_c \ \1 $                                   &  $c_1 =  $  & $481 \pm5$  & $482\pm5$ &  $484\pm4$ &  $484\pm4$ & $498\pm3$\\
$O_2 = \frac{1}{N_c}\ell^i S^i$                     &  $c_2 =  $  & $-47 \pm39$ & $-30\pm34$ & $-31\pm20$ & $8\pm15$ & $38\pm34$\\
$O_3 = \frac{1}{N_c}S^iS^i$                         &  $c_3 =  $  & $161\pm 16$ & $149\pm11$ & $159\pm16$ & $149\pm11$ & $156\pm16$\\
$O_4 = \frac{1}{N_c}T^aT^a$                         &  $c_4 =  $  & $169\pm36$  & $170\pm36$ & $138\pm27$ & $142\pm27$ &\\
$O_5 = \frac{3}{N_c^2}\ell^{(2)ij}G^{ia}G^{ja}$     &  $c_5 =  $  & $-443\pm459$&            & $-371\pm456$&        & $-514\pm458$\\
$O_6 = \frac{1}{N_c^2}\ell^iT^aG^{ia}$              &  $c_6 =  $  & $473\pm355$ & $433\pm353$ &            &      & $-606\pm273$
\vspace{0.15cm} \\
\hline
$\chi_{\mathrm{dof}}^2$                             &             & $0.43$      & $0.68$ & $1.1$           & $0.96$ & $11.5$\\
\hline 
\end{tabular}}
\end{table}

The states belonging to $[{\bf 70}, \ell]$-plets are more 
difficult due to the presence of a mixed symmetry.  So far, in calculating 
the mass spectrum, the general practice 
was to split the baryon into an excited quark and a symmetric core,
in this way reducing  the problem to the well known ground state.
There are two drawbacks in this procedure.
One is that each generator of SU(2$N_f$) is written as a sum of two
terms, one acting on the excited quark and the other on the core. 
As a consequence, 
the number of linearly
independent operators to be used in Eq. (\ref{OLFS})
increases tremendously  and the number of coefficients 
to be determined becomes larger or much larger than the experimental data
available. For example, for the $[{\bf 70},1^-]$ multiplet with $N_f = 2$ 
one has 12 linearly independent
operators up to order $1/N_c$  \cite{CCGL}, instead of 6 
as in the present approach (see Table \ref{operators}). 
We recall that there are only 7 nonstrange 
resonances belonging to this band. They are given in Table \ref{MASSES}.
Consequently, in selecting the most dominant operators one has to make an
arbitrary choice  \cite{CCGL}.
The second drawback is  due to the 
truncation of  the wave function containing the
orbital and the spin-flavor parts. The exact wave function is
given by a linear combination of terms of equal weight where each term 
corresponds to a given Young tableau of mixed symmetry denoted 
by $[N_c-1,1]$.
In the above procedure
only the term where the last quark is in the second row
was kept, the other $N_c - 2$ terms being ignored. 
As a consequence the order of the spin-orbit operator became $N^0_c$.

In this practice
the matrix elements of the excited quark are straightforward, as being
described by single-particle operators. The matrix elements of the core operators 
$S^i_c$ and $T^a_c$ are also simple to calculate, while  $G^{ia}_c$
are more involved.
Analytic group theoretical formulas for several types of matrix elements of 
the SU(4) generators  have been derived in the late sixtieths  \cite{HP},
in the context of nuclear physics. 
Every matrix element is factorized 
according to a generalized Wigner-Eckart theorem into a reduced
matrix element and an SU(4) Clebsch-Gordan coefficient. 
Recently we have extended the approach to calculate isoscalar factors
needed for the matrix elements of
SU(6) generators between symmetric $[N_c]$ states\  \cite{Matagne:2006xx}.

\vspace{0.3cm}
{\it The $[\bf 70,1^-]$-plet.}
Here we propose a new method where the splitting into an excited quark and
a core is unnecessary. Details can be found in Ref. \ \cite{Matagne:2006dj}. 
All one needs to know are the matrix
elements of the SU(2$N_f$) generators between mixed symmetric
states   $[N_c-1,1]$. For $N_f$ = 2 these are provided by the 
work of Hecht and Pang \cite{HP}.
To our knowledge such matrix elements are yet 
unknown for  $N_f$ = 3. Thus our work deals with nonstrange 
baryon resonances only. 
The list of operators $O_i$ contributing to 
the mass operator (\ref{massoperator}) 
is shown in Table \ref{operators}
together with the coefficients $c_i$. The first column gives 
all linearly independent operators of type (\ref{OLFS}) up to order $1/N_c$.
The other columns indicate the values obtained for the dynamical 
coefficients $c_i$ from various fits. Fit 1 contains all operators and gives 
the best $\chi_{\mathrm{dof}}^2$. In the other columns one or two operators 
have been removed in order to understand their role in the fit. One can see 
that the removal of $O_5$ or of $O_6$ is not so dramatic but that
the removal of $O_4$ badly deteriorates the fit. The data are from 
Ref. \  \cite{Yao}.
It is important to note that here the angular momentum operator of components
${\ell_i}$ ($i$ = 1,2,3) is an intrinsic operator, acting on the entire 
system, not  on the $N_c$-th quark only, as in previous studies,
which means it was taken relative to a fixed center of mass
\ \cite{CCGL}. In other words here we do not have a center of mass problem
which may complicate the $N_c$ counting. 
\begin{center}
\begin{table}[h!]
\caption{Partial contributions and total mass (MeV) 
of $[\bf 70,1^-]$ resonances predicted by the 
$1/N_c$ expansion. 
The last two columns reproduce the experimental  masses and the status 
of resonances \cite{Yao}.}
\label{MASSES}
\renewcommand{\arraystretch}{2.5}
{\tiny
\begin{tabular}{crrrrrrcccl}\hline 
                    &      \multicolumn{6}{c}{Part. contrib. (MeV)}  & \hspace{.0cm} Total (MeV)   & \hspace{.0cm}  Exp. (MeV)\hspace{0.0cm}& &\hspace{0.cm}  Name, status \hspace{.0cm} \\

\cline{2-7}
                    &   \hspace{.0cm}   $c_1O_1$  & \hspace{.0cm}  $c_2O_2$ & \hspace{.0cm}$c_3O_3$ &\hspace{.0cm}  $c_4O_4$ &\hspace{.0cm}  $c_5O_5$ & $c_6O_6$   &        \\
\hline
$^2N_{\frac{1}{2}}$        & 1444 & -16 & 40 & 42 & 0 & -13  &   $1529\pm 11$  & $1538\pm18$ & & $S_{11}(1535)$****  \\
$^4N_{\frac{1}{2}}$        & 1444 &  39 & 201& 42 & -31& -33 &   $1663\pm 20$  & $1660\pm20$ & & $S_{11}(1650)$**** \\
$^2N_{\frac{3}{2}}$        & 1444 & -8  & 40 & 42 & 0  &  7  &   $1525\pm 8$   & $1523\pm8$  & & $D_{13}(1520)$****\\
$^4N_{\frac{3}{2}}$        & 1444 & 16  & 201& 42 & 25 & -13 &   $1714\pm45$   & $1700\pm50$ & & $D_{13}(1700)$***\\
$^4N_{\frac{5}{2}}$        & 1444 & -24 & 201& 42  & -6 & 20 &   $1677\pm8$    & $1678\pm8$  & & $D_{15}(1675)$****\\
\hline
$^2\Delta_{\frac{1}{2}}$  &  1444 & 16  & 40 & 211 & 0  & -66   & $1645\pm30$  & $1645\pm30$ & & $S_{31}(1620)$**** \\
$^2\Delta_{\frac{3}{2}}$  &  1444 & -8  & 40 & 211 & 0  & -33   & $1720\pm50$  & $1720\pm50$ & & $D_{33}(1700)$**** \vspace{0.15cm} \\
\hline 
\end{tabular}}
\end{table}
\end{center}

The partial contributions and the total mass predicted by the 
$1/N_c$ expansion are given in Table 2.
One can see that the contributions of all terms containing
angular momentum, and in particular that of the spin-orbit,   
are small. This can be viewed as a dynamical effect. An entirely new 
quantitative result is that the isospin-isospin 
term $O_4$ brings a dominant contribution to $\Delta$ resonances, of the
same order as the spin-spin term $O_3$ brings to $N$ resonances.  

\vspace{0.3cm}
{\it Conclusions.} 
These results shed a new light into the description 
of the baryon multiplet  $[{\bf 70},1^-]$ in the  $1/N_c$ expansion. 
The main findings are:
\begin{itemize}
 \item In the mass formula the expansion starts at order $1/N_c$, as for the ground 
 state, instead of $N^0_c$ as previously concluded.
 \item The isospin operator  $O_4 = \frac{1}{N_c} T^aT^a$ is crucial in the
 fit to the existing data and its contribution is as important as that 
 of the spin term $O_3 = \frac{1}{N_c} S^iS^i$. 
 \end{itemize}
 It would be interesting to reconsider the study of higher excited 
  baryons, for example those belonging to $[{\bf 70},\ell^+]$ multiplets, 
  in the spirit of the present approach.
  Based on group theoretical  arguments it is expected that the mass 
  splitting starts at order $1/N_c$, as a general rule.


\end{document}